OpenStudy (anonymous):

Find the indefinite integral of cos(x^1/2) dx ATTEMPT: I let u = x^1/2 Then, du = 1/2(x^-1/2)dx <-- but that term is not in the integrand..

OpenStudy (anonymous):

Nevermind, I remember.

OpenStudy (phi):

in case someone is wondering: \[ du= \frac{1}{2\sqrt{x} }dx \] \[ dx= 2 u du \] and the integral is \[\int\limits u \cos(u) du\] which tackled using integration by parts